OpenStudy (anonymous):

Suppose csc θ = √5 Find the exact value of cos 2θ

OpenStudy (lgbasallote):

note that \(\csc \theta = \frac{1}{\cos \theta}\)

OpenStudy (anonymous):

 can you take it from here?